OpenStudy (anonymous):

evaluate 4 / (sqrt (1+2x))^3

OpenStudy (anonymous):

i don't know whether the answer is \[\frac{ -8 }{ \sqrt{1+2x} } +C\] or -\[\frac{ -4 }{ \sqrt{2x+1} } +C\]

OpenStudy (anonymous):

evaluate \[ \int\limits_{}^{}\frac{ 4 }{ (\sqrt{1+2x})^{3} } dx\]

OpenStudy (kc_kennylau):

http://prntscr.com/29m3p6

OpenStudy (anonymous):

its not \[\int\limits\limits_{}^{} \frac{ 4 }{ (\sqrt{1+2x})^{3/2} } dx\]

OpenStudy (anonymous):

its \[\int\limits\limits_{}^{} \frac{ 4 }{ (\sqrt{1+2x})^{3} } dx\]

OpenStudy (kc_kennylau):

\[\Large(\sqrt{1+2x})^3=(1+2x)^{\frac32}\]

OpenStudy (kc_kennylau):

Welcome to OS anyway :) Hope you have a good time here :)

OpenStudy (anonymous):

i'm sorry!! :( i so careless

OpenStudy (kc_kennylau):

not your fault :)

OpenStudy (anonymous):

thanks for ur warm welcome. :)

OpenStudy (kc_kennylau):

:)